OpenStudy (anonymous):

3x^2-5x=2 solve the 2 quad equations 2x^2-15x=8

OpenStudy (anonymous):

i've just lost my answer what happened? try again:

OpenStudy (anonymous):

3x^2 - 5x - 2 = 0 (3x +1)(x - 2) = 0 x = -1/3, 2 2x^2-15x - 8 = 0 (2x + 1 )(x - 8 ) = 0 s = -1/2, 8

OpenStudy (anonymous):

s = x lol

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

yes